# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v4

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can guide the reconstruction of an undersampled subsequent contrast. To this end, several end-to-end learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach addressing this issue, additionally providing an explanatory framework for the multi-contrast problem based on the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Consequently, incorporating prior information into the reconstruction reduces to a simple replacement of the aliased content of the reconstruction iterate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-CoSMo. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the public NYU fastMRI DICOM dataset, showing improved generalizability compared to end-to-end methods, and on two in-house multi-coil raw datasets, offering up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM.
@@ 1 Introduction
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan (Ehrhardt and Betcke, 2016; Weizman et al., 2016; Zhou and Zhou, 2020) and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously (Bilgic et al., 2011; Huang et al., 2014; Kopanoglu et al., 2020). In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
@@ 2.1 Reconstruction methods for accelerated MRI
Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) (Lustig et al., 2007), which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA (Chambolle et al., 1998) or ADMM ((Boyd et al., 2011)) family of algorithms. Most modern deep learning-based reconstruction methods focus on improving the denoising part. Plug-and-play (PnP) methods (Ahmad et al., 2020; Kamilov et al., 2023) replace the citep operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model. Unrolled networks (Schlemper et al., 2017; Aggarwal et al., 2018; Pezzotti et al., 2020) extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end. This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability (Ahmad et al., 2020).
@@ 2.2 Unpaired image-to-image modeling
MUNIT (Huang et al., 2018) extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT. The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same ”content” but differing in ”style”. More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains. And unlike other content/style modeling frameworks such as (Kotovenko et al., 2019; Kwon and Ye, 2021; Wu et al., 2023), MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. Compared to Denck et al. (2021), we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast. We model using an approach based on the MUNIT framework, referring to the shared and non-shared components as content and style, respectively.
@@ 2.3 Combining image translation with reconstruction
Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of joint image synthesis and reconstruction, e.g. PROSIT (Mattern et al., 2020) and rsGAN (Dar et al., 2020), involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction. More recent work by Xuan et al. (2022) proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory (Wang et al., 2024). Levac et al. (2023) formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style. This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure that is available to be learned in this data and utilized in the reconstruction task.
@@ 3.2 Content/Style modeling
Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images. We formulate our content/style model based on the MUNIT framework (Huang et al., 2018).
@@ 3.2.1 Unpaired MUNIT pre-training
We define an image domain as the set of images of a certain contrast comprising the dataset, where . ”Content” is defined as the underlying contrast-independent structure and is represented as a set of feature maps, whereas ”style” corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector. MUNIT posits the existence of functions and their inverses , and learns them jointly via unpaired training, given samples from marginal distributions . In practice, the encoder is split into content encoder and style encoder . Thus, the content/style model is specified as .
@@ 4.1 Experiment design
We conducted four sets of experiments to comprehensively evaluate PnP-CoSMo – (a) analyses on simulated MRI images, (b) benchmark against end-to-end methods on the public NYU fastMRI DICOM data, (c) benchmark and ablation on clinical multi-coil raw data, and (d) a preliminary radiological evaluation. Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications, except in the NYU benchmark where the reverse case of T2W-guided T1W reconstruction was considered as well.
@@ 4.1.1 Analyses on simulated MR data
The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings. To this end, we used simulated T1W and T2W images based on BrainWeb phantoms (Collins et al., 1998). BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split into the model training (18), model validation (1), and reconstruction testing (1) subsets. T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges. For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations. We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-CoSMo algorithm, and the effectiveness, robustness, and tuning of the CR module.
@@ 4.1.2 Benchmark on public NYU fastMRI DICOM data
In the NYU benchmark, we compared PnP-CoSMo against end-to-end reconstruction methods. The question we sought to answer was whether or not PnP-CoSMo, which requires only image-domain training data, can outperform end-to-end methods that additionally require k-space training data and are optimized to directly solve the reconstruction problem. Furthermore, end-to-end methods are task-specific models where the reference and target contrasts are fixed, e.g. T1W as reference and T2W as target. On the other hand, the content/style model used in PnP-CoSMo does not possess this kind of uni-directionality. Hence, we sought to assess the generalizability of PnP-CoSMo across two different reconstruction tasks, namely T1W-guided T2W reconstruction and T2W-guided T1W reconstruction.
@@
The public NYU fastMRI brain dataset (Knoll et al., 2020; Zbontar et al., 2018) includes multi-coil raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. For multi-contrast experiments, however, the NYU raw dataset does not provide subject labels for pairing the contrasts. Therefore, instead of the NYU raw data, we used the NYU DICOMs in this study, following Xuan et al. (2022). A paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil k-space data was simulated via Fourier transform with 1D Cartesian random sampling at and added Gaussian noise of . The dataset was split into 3 folds at the subject-level for cross-validation, with two partitions representing training data and one representing the test data in each fold. For PnP-CoSMo, images from the training set were used to pre-train and fine-tune the content/style model, whereas the end-to-end methods were trained on paired examples composed of the reference image, the target contrast k-space, and its ground truth image. We benchmarked PnP-CoSMo against one well-known single-contrast network – MoDL (Aggarwal et al., 2018) – and two recent multi-contrast networks – MTrans (Feng et al., 2022) and MC-VarNet (Lei et al., 2023). These networks were evaluated on the aforementioned two reconstruction tasks, assessing both their in-distribution and out-of-distribution performance – e.g. for the T1W-guided T2W reconstruction task, in-distribution means that the end-to-end model was trained on this task, whereas out-of-distribution means the model was trained of the T2W-guided T1W reconstruction task.
@@
| LUMC-TRA | model-trainI | 295 | 418 | 360 / 415 |
| model-valI | 16 | 17 | 17 / 17 | |
@@
| LUMC-COR | model-trainI | 242 | 277 | 269 / 272 |
| model-valI | 18 | 18 | 18 / 18 | |
@@ 4.1.3 Benchmark and ablation on in-house LUMC multi-coil datasets
Each of the two LUMC datasets was split into model-train, model-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate the content/style models. The models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of =1.8-2. We further undersampled it retrospectively to higher accelerations by dropping subsets of the acquired lines. An overview of the data split is shown in Table 2. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.

### v5

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can guide the reconstruction of an undersampled subsequent contrast. To this end, several end-to-end learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach that does not require any k-space training data, relying solely on image-domain datasets, a large part of which can be unpaired. Additionally, our approach provides an explanatory framework for the multi-contrast problem based on the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Consequently, incorporating prior information into the reconstruction reduces to a simple replacement of the aliased content of the reconstruction iterate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-CoSMo. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the public NYU fastMRI DICOM dataset, showing improved generalizability compared to end-to-end methods, and on two in-house multi-coil raw datasets, offering up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM.
@@ 1 Introduction
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types: (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan (Ehrhardt and Betcke, 2016; Weizman et al., 2016; Zhou and Zhou, 2020) and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously (Bilgic et al., 2011; Huang et al., 2014; Kopanoglu et al., 2020). In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
@@ 2.1 Reconstruction methods for accelerated MRI
Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) (Lustig et al., 2007), which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA (Chambolle et al., 1998) or ADMM ((Boyd et al., 2011)) family of algorithms. Most modern deep learning-based reconstruction methods focus on improving the denoising part. Plug-and-play (PnP) methods (Ahmad et al., 2020; Kamilov et al., 2023) replace the proximal operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model. Unrolled networks (Schlemper et al., 2017; Aggarwal et al., 2018; Pezzotti et al., 2020) extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end. This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability (Ahmad et al., 2020).
@@ 2.2 Unpaired image-to-image modeling
MUNIT (Huang et al., 2018) extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT. The result was a stochastic image translation model, which, given an input image, generates a distribution of synthetic images sharing the same ”content” but differing in ”style”. More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains. And unlike other content/style modeling frameworks such as (Kotovenko et al., 2019; Kwon and Ye, 2021; Wu et al., 2023), MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. Compared to Denck et al. (2021), we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast. We model using an approach based on the MUNIT framework, referring to the shared and non-shared components as content and style, respectively.
@@ 2.3 Combining image translation with reconstruction
Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of joint image synthesis and reconstruction, e.g. PROSIT (Mattern et al., 2020) and rsGAN (Dar et al., 2020), involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction. More recent work by Xuan et al. (2022) proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory (Wang et al., 2024). Levac et al. (2023) (whose approach we refer to as ”PnP-Diffusion”) formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style. This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure that is available to be learned in this data and utilized in the reconstruction task.
@@ 3.2 Content/Style modeling
Given an image dataset of two MR contrasts, we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images. We formulate our content/style model based on the MUNIT framework (Huang et al., 2018).
@@ 3.2.1 Unpaired MUNIT pre-training
We define an image domain as the set of images of a certain contrast comprising the dataset, where . ”Content” is defined as the underlying contrast-independent structure and is represented as a set of feature maps, whereas ”style” corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector.
MUNIT posits the existence of functions and their inverses , and learns them jointly via unpaired training, given samples from marginal distributions . In practice, the encoder is split into content encoder and style encoder . Thus, the content/style model is specified as .
@@ 4.1 Experiment design
We conducted four sets of experiments to comprehensively evaluate PnP-CoSMo – (a) analyses on simulated MRI images, (b) benchmark against end-to-end methods on the public NYU fastMRI brain data (Knoll et al., 2020; Zbontar et al., 2018), (c) benchmark and ablation on clinical multi-coil raw data, and (d) a preliminary radiological evaluation. Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications, except in the NYU benchmark where the reverse case of T2W-guided T1W reconstruction was considered as well.
@@ 4.1.1 Analyses on simulated MR data
The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings. To this end, we used simulated T1W and T2W images based on BrainWeb phantoms (Collins et al., 1998). BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split into the model training (18), model validation (1), and reconstruction testing (1) subsets. 2D T1W and T2W spin-echo scans in the transversal plane were simulated using TE/TR values randomly sampled from realistic ranges, specifically TR and TE for T1W scans, and TR and TE for T2W scans. For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations. We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-CoSMo algorithm, and the effectiveness, robustness, and tuning of the CR module.
@@ 4.1.2 Benchmark on public NYU fastMRI DICOM data
In the NYU benchmark, we compared PnP-CoSMo against state-of-the-art reconstruction methods belonging to both end-to-end and plug-and-play paradigms. First, we sought to answer the following question – can PnP-CoSMo, which requires only image-domain training data, outperform end-to-end methods ,which additionally require k-space training data and are optimized to directly solve the reconstruction problem? Furthermore, end-to-end methods are task-specific models where the reference and target contrasts are fixed, e.g. T1W as reference and T2W as target. On the other hand, the content/style model used in PnP-CoSMo does not possess this kind of uni-directionality. Hence, we sought to assess the generalizability of PnP-CoSMo across two different reconstruction tasks: (a) T2W reconstruction with guidance from T1W reference and (b) T1W reconstruction with guidance from T2W reference. Finally, we compared PnP-CoSMo against other plug-and-play methods, which share a similar design pattern as ours where the modeling/learning problem is decoupled from the reconstruction problem and the requirement is only image-domain training data.
@@
The public NYU fastMRI brain dataset includes multi-coil raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. For multi-contrast experiments, however, the NYU raw dataset does not provide subject labels for pairing the contrasts. Therefore, instead of the NYU raw data, we used the NYU DICOMs in this study, following Xuan et al. (2022). A paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans, all of them being either 2D spin echo or turbo spin echo acquisitions in the transverse plane with median in-plane FOV of 220220 mm2 for both, in-plane voxel size of 0.690.69 mm2 for T1W scans and 0.570.57 mm2 for T2W scans, slice thickness of 5 mm for both, FA=20-150∘, TR=220-874 ms, TE=2.46-17 ms, and ETL=1-5 for T1W scans, and FA=90-180∘, TR=2696-8910 ms, TE=9.9-119 ms, and ETL=11-27 for T2W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil k-space data was simulated via Fourier transform with 1D Cartesian random sampling at and added Gaussian noise of .
The dataset was split into three folds at the subject-level for cross-validation based on random shuffling, with two partitions representing training data and one representing the test data in each fold. For PnP-CoSMo, all images from the training set were used to pre-train (ignoring the pairing information) and fine-tune (utilizing the pairing information) the content/style model, whereas the end-to-end methods were trained on paired examples composed of the reference image, the target contrast k-space, and its ground truth image. The plug-and-play baselines, were trained, like PnP-CoSMo, on the full image-domain data in the cross-validation training folds.
We benchmarked PnP-CoSMo against three end-to-end methods and three plug-and-play methods. Among the end-to-end baselines we included one well-known single-contrast network, namely MoDL (Aggarwal et al., 2018), and two recent multi-contrast networks, namely MTrans (Feng et al., 2022) and MC-VarNet (Lei et al., 2023). These networks were evaluated on the aforementioned two reconstruction tasks, assessing both their in-distribution and out-of-distribution performance. Specifically, for the T2W reconstruction task, ”in-distribution” means that the end-to-end model was trained to reconstruct T2W images, whereas ”out-of-distribution” means the model was trained to reconstruct T1W images but applied on the T2W reconstruction task, and vice versa. While MC-VarNet and MTrans used the reference image information, MoDL ignored the reference information. Among the plug-and-play baselines we included a single-contrast method, namely PnP-CNN (Ahmad et al., 2020), and two guided methods, namely PnP-Diffusion (Levac et al., 2023) and PROSIT (Mattern et al., 2020). Note that these methods, like PnP-CoSMo, are not contrast-task-specific in the same way end-to-end models are, and thus the concept of in-distribution and out-of-distribution tasks does not apply to them.
@@
| LUMC-TRA | image-trainI | 295 | 418 | 360 / 415 |
| image-valI | 16 | 17 | 17 / 17 | |
@@
| LUMC-COR | image-trainI | 242 | 277 | 269 / 272 |
| image-valI | 18 | 18 | 18 / 18 | |
@@ 4.1.3 Benchmark and ablation on in-house LUMC multi-coil datasets
Each of the two LUMC datasets was split into image-train, image-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate the content/style models. The models were pre-trained on the full image-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the image-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of =1.8-2. We further undersampled it retrospectively to higher accelerations by dropping subsets of the acquired lines. The recon-val split was used to tune the reconstruction algorithm hyperparameters (such as CR step size for PnP-CoSMo, regularization weights for classical baselines, etc.), whereas recon-test was the held-out data on which the final metrics are reported. An overview of the data split is shown in Table 2. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.